OpenStudy (anonymous):

When the quadratic equation y = (x − 5)(2x + 3) is written in standard form, what is the value of the coefficient "b"?

myininaya (myininaya):

y=(x-5)(2x+3)=x(2x)+x(3)-5(2x)-5(3) =2x^2+3x-10x-15 =2x^2-7x-15 =ax^2+bx+c b=-7